Imperial Jewelers manufactures and sells a gold bracelet for $403.00. The company’s accounting system says that the unit product cost for this bracelet is $267.00 as shown below:

Direct materials $ 143
Direct labor 85
Manufacturing overhead 39
Unit product cost $ 267

The members of a wedding party have approached Imperial Jewelers about buying 16 of these gold bracelets for the discounted price of $363.00 each. The members of the wedding party would like special filigree applied to the bracelets that would require Imperial Jewelers to buy a special tool for $454 and that would increase the direct materials cost per bracelet by $8. The special tool would have no other use once the special order is completed.

To analyze this special order opportunity, Imperial Jewelers has determined that most of its manufacturing overhead is fixed and unaffected by variations in how much jewelry is produced in any given period. However, $9.00 of the overhead is variable with respect to the number of bracelets produced. The company also believes that accepting this order would have no effect on its ability to produce and sell jewelry to other customers. Furthermore, the company could fulfill the wedding party’s order using its existing manufacturing capacity.

Required:

1. What is the financial advantage (disadvantage) of accepting the special order from the wedding party?

Financial advantage I   
Answer #1

Revenue from special order = 16*363 = 5808

Cost of making

= (143+8+85+9)*16 + 454 = 4374

Financial advantage = 5808 - 4374

Financial advantage = 1434

Comment if you face any issues